This is a word problem my teacher made up for us. I am having trouble setting it up. Please help! "The difference of two numbers is 5. Twice the larger number is three times the smaller number. Find the two numbers."
Answer by stanbon(48558)

"The difference of two numbers is 5. Twice the larger number is three times the smaller number. Find the two numbers."
---------
Let the two numbers be x and x+5 so the difference will be 5.
EQUATION:
Assume (x+5) is the larger number so you get.
2(x+5)=3x
2x+10=3x
x=10 (this is the smaller number)
x+5=15 (this is the larger number)
Cheers,
Stan H.